Coastal Computer operates two retail outlets in Oakview, one on Main Street and the other in Lakeland Mall. The stores share the use of a central accounting department. The cost of the accounting department for last year was $208,000. Following are the operating results for the two stores for the year ain StreetL Lakeland Mall Sales revenues Number of computers sold s 1,913,470 2,655 2,753 530 1,845 Required: a. Allocate the cost of the central accounting department to the two stores based on 1. Number of computers sold. (Do not round intermediate calculations.) Main Street Lakeland Mall Total Allocated accounting department cost 2. Store revenue. (Do not round intermediate calculations.) Main Street Lakeland Mall Total Allocated accounting department cost
Answer #1

Answer

  • All working forms part of the answer
  • Requirement ‘a’ 1

Note: Total computer sold = 2655 + 1845 = 4,500

Main Street

Lakeland Mall

Total

Allocated accounting department cost

$122,720

$85,280

$208,000

Working

[$ 208000 x (2655 computers/4500 computers)]

[$ 208000 x (1845 computers/4500 computers)]

  • Requirement ‘a’ 2

Note: Total Revenue = 1913470 + 2753530 = $ 4667000

Main Street

Lakeland Mall

Total

Allocated accounting department cost

$85,280

$122,720

$208,000

Working

[$ 208000 x ($1913470 revenue/4667000 total revenue)]

[$ 208000 x ($2753530 revenue/4667000 total revenue)]